The induced electric field 13 cm from the axis of a solenoid with a 9.0 cm radius is 44 V/m.

Find the rate of change of the solenoid's magnetic field.
Answer #1
Concepts and reason

The concepts required solve the given problem are Faraday’s law of induction, and definition of magnetic flux.

Initially, using the Faraday’s law and magnetic flux expression find out the equation for the rate of change of magnetic field.

Then find the induced emf{\rm{emf}} in the solenoid by using expression for induced emf{\rm{emf}} in terms of electric field.

Fundamentals

The magnetic flux through a closed loop of current carrying wire can be given by the following expression:

ϕ=B.A\phi = \vec B.\vec A

Here, B is the magnetic field and A is the area of the loop.

The expression of the emf generated within a solenoid due to the change in magnetic flux passing through it is as follows:

ε=dϕdt\varepsilon = - \frac{{d\phi }}{{dt}}

Here, dϕdt\frac{{d\phi }}{{dt}} is the rate of change of magnetic flux through the solenoid with respect to time.

The expression of the induced emf{\rm{emf}} in terms of electric field is give as follows:

ε=Edl\varepsilon = - \int {E \cdot dl}

According to Faraday’s law, the induced emf{\rm{emf}} is given as follows:

ε=dϕdt\varepsilon = - \frac{{d\phi }}{{dt}}

Substitute BABA for ϕ\phi in the above equation.

ε=d(BA)dt=AdBdt\begin{array}{c}\\\varepsilon = - \frac{{d\left( {BA} \right)}}{{dt}}\\\\ = - A\frac{{dB}}{{dt}}\\\end{array}

Rearrange the above equation for change in magnetic field.

dBdt=εA\frac{{dB}}{{dt}} = - \frac{\varepsilon }{A}

The emf{\rm{emf}} induced in terms of electric field is give as follows:

ε=Edl=Edl\begin{array}{c}\\\varepsilon = - \int {E \cdot dl} \\\\ = - E\int {dl} \\\end{array}

The integral dl\int {dl} value is equal to 2πr.2\pi r.

Substitute 2πr2\pi r for dl\int {dl} in the above equation.

ε=E(2πr)=2Eπr\begin{array}{c}\\\varepsilon = - E\left( {2\pi r} \right)\\\\ = - 2E\pi r\\\end{array}

Substitute 2Eπr - 2E\pi r for ε\varepsilon in the above equation dBdt=εA.\frac{{dB}}{{dt}} = - \frac{\varepsilon }{A}.

dBdt=(2Eπr)A=2EπrA\begin{array}{c}\\\frac{{dB}}{{dt}} = - \frac{{\left( { - 2E\pi r} \right)}}{A}\\\\ = \frac{{2E\pi r}}{A}\\\end{array}

The cross section area A is give as followed:

A=πR2A = \pi {R^2}

Substitute πR2\pi {R^2} for A in the above equation.

dBdt=2EπrπR2=2ErR2\begin{array}{c}\\\frac{{dB}}{{dt}} = \frac{{2E\pi r}}{{\pi {R^2}}}\\\\ = \frac{{2Er}}{{{R^2}}}\\\end{array}

Substitute 44 V/m for E, 9.0 cm for R, and 13 cm for r in the above equation.

dBdt=2(44V/m)(13cm)(1m100cm)(9.0cm)2(1m2104cm2)=1412T/s\begin{array}{c}\\\frac{{dB}}{{dt}} = \frac{{2\left( {44\,{\rm{V/m}}} \right)\left( {13\;{\rm{cm}}} \right)\left( {\frac{{1\,{\rm{m}}}}{{100\,{\rm{cm}}}}} \right)}}{{{{\left( {9.0\,{\rm{cm}}} \right)}^2}\left( {\frac{{1\;{{\rm{m}}^2}}}{{{{10}^4}\,{\rm{c}}{{\rm{m}}^2}}}} \right)}}\\\\ = 1412\,{\rm{T/s}}\\\end{array}

Ans:

The rate of change of solenoid’s magnetic field is 1412 T/s.